# Geometrical Aspects Of Resources Distribution In Quantum Random Circuits

### v2

@@ Abstract
In this work, we explore how resources are distributed among the states generated by quantum random circuits (QRC). We focus on multipartite non-locality, but we also analyze quantum correlations by appealing to different entanglement and non-classicality measures. We compare universal vs non-universal sets of gates to gain insight into the problem of explaining quantum advantage. By comparing the results obtained with ideal (noiseless) vs noisy intermediate-scale quantum (NISQ) devices, we lay the basis of a certification protocol, which aims to quantify how robust is the resources distribution among the states that a given device can generate.
keywords:
Quantum random circuits $\cdot$ Quantum non-locality $\cdot$ Quantum entanglement $\cdot$ CHSH-Mermin-Svetlichny inequalities
[…]
[inst1]organization=University of Cagliari, addressline=Via Is Mirrions 1, city=Cagliari, postcode=09123, state=Italy
[…]
[inst2]organization=Universidad Nacional de La Plata, Instituto de Fisica (IFLP-CCT-CONICET), addressline=C.C. 727, city=La Plata, postcode=1900, state=Argentine
[…]
[inst3]organization=Technische Universität München. Institute for Advanced Study (IAS), addressline=Lichtenbergstraße 2 a, city=München, postcode=80333, state=Germany
1 Introduction
@@
Quantum computers offer promising computational advantages for the years to come. There exist many operating devices nowadays, which can be easily accessed using the cloud quantum computing model. This rapid growth is accompanied by claims of quantum advantage with near term devices [1, 2]. Though the interpretation of the obtained results generates a heated debate and should be taken cautiously [3], it is clear that the road map is promising. There are great chances that, in the near future, quantum computers will be able to perform tasks that are very hard for classical supercomputers, even if the path to fault tolerant universal quantum computing remains challenging.
@@
There are many hypotheses to explain why quantum computers seem to have a speed-up with regard to their classical cousins. Overall, the reasons are not completely clear. Recent studies point to different resources, such as entanglement and non-locality [4], coherence [5], and contextuality [6]. But none of these quantum features seem to suffice on its own to explain quantum advantage, due to results such as the Gottesmann-Knill theorem [7] (though see for example the discussion in [8]). Furthermore, the certification of a quantum device is usually a complicated business. As extant quantum processors are very sensitive to noise, it is important to develop different tools to assess their capabilities (see for example [9, 10, 11, 12]). The general question could be posed in this way: how are quantum resources distributed among the states that a concrete quantum device is able to produce?
@@
Here we study the resources produced in different architectures of quantum computers using quantum random circuits (QRC) [13]. We focus on non-locality by using the Mermin [14] and the Svetlichny inequalities [15, 16]. We also study entanglement by appealing to different measures, such as tangle and concurrence (see A of this work). We first perform a theoretical study based on numerical simulations using Qiskit [17] and Amazon Braket SDK [18]. We analyze the performance of the devices with different levels of noise and number of shots, and compare universal (Clifford $+$ T) vs non-universal (Clifford) sets of elementary gates. Given a set of physical features which are thought of as a resource for quantum computing (as for example, non-locality), we quantify to which degree a device can generate a state space rich enough to generate the desired resources.
The paper is organized as follows. In Section 2 after giving a very brief overview of QRC, we show the results of the simulations of violation of Mermin and Svetlichny inequalities for different qubits numbers. In Section 3, we present results regarding entanglement for comparison purposes. Finally, in Section 6, we draw our conclusions.
2 Quantum Random Circuits: distribution of resources
QRC are used in different areas of quantum information science [13]. In order to give a general idea, consider a scheme of circuit generation as depicted in Figure 1. On each stage of the protocol, random unitary gates are chosen and applied to the circuit. As a result, a random circuit is obtained. Remarkably, QRC are a promising tool to show quantum advantage [1]. Here, we use them to show how quantum resources are distributed among the possible states that a quantum computer is able to generate. An ideal quantum computer should be able to reach all possible states of a system of $N$-qubits. A real quantum device is able to implement a given set of elementary quantum gates on a fixed set of qubits. By generating QRC with different depths, sets of elementary gates, and levels of noise, we study how robust is the resources distribution with regard to the ideal case.
Specifically, by comparing Clifford (H, S and CNOT) vs non Clifford (Clifford $+$ T) sets of gates, we highlight the differences of how quantum resources are distributed between universal vs non-universal quantum computing.

### v3

# Distribution of Non-Locality On Quantum Random Circuits

@@ Abstract
In this work we explore how different types of resources are distributed among the states generated by quantum random circuits (QRC). We focus on multipartite non-locality, but we also analyze quantum correlations by appealing to different entanglement and non-classicality measures. We analyze the violation of Mermin and Svetlichny inequalities in order to get a glance at the distribution of nonlocality and genuine multipartite nonlocality. Next, we compare universal vs non-universal sets of gates, to gain insight into the problem of explaining quantum advantage. By comparing the results obtained with ideal (noiseless) vs noisy intermediate-scale quantum (NISQ) devices, we lay the basis of a certification protocol, which aims to quantify how robust is the resources distribution among the states that a given device can generate. We have implemented our non-locality-based benchmark on actual quantum processors with different architectures, in order to assess up to which point they are capable of reproducing the ideal results.
preprint: APS/123-QED
I Introduction
@@
Quantum computers offer promising computational advantages for the years to come. There exist many operating devices nowadays that can be easily accessed using the cloud quantum computing model. This rapid growth is accompanied by claims of quantum advantage with near-term devices [1, 2, 3]. Though the interpretation of quantum supremacy claims generates a heated debate and should be taken cautiously (see for example [4] and [5]), it is clear that the road map is promising. There are great chances that, in the near future, quantum computers will be able to perform tasks that are very hard for classical supercomputers, even if the path to fault-tolerant universal quantum computing and commercial applications remains challenging.
Nowadays prototypes of quantum computers are noisy and present a variable performance among the different architectures. It is then relevant to characterize how capable these devices are of producing different types of resources that are considered nonclassical and necessary for computational advantage. In this work we focus on the capability of simulating nonclassical correlations using quantum random circuits. Our study shows how rich is the quantum state space a given quantum device can produce. We also analyze the differences of devices using only Clifford vs Clifford ++ T sets of native gates. As is well known, the Clifford ++ T set can be used to approximate any conceivable quantum circuit, given that it is a universal set. On the contrary, a QPU using only Clifford gates is not universal, and it is known that its outputs can be efficiently simulated by a classical probabilistic Turing machine [6]. Out of our comparison, we obtain valuable information about how resources are distributed in both cases, yielding a novel insight on discussions about quantum advantage.
@@
There are many hypotheses to explain why quantum computers seem to have a speed up with regard to their classical cousins. Overall, the reasons are not completely clear. Recent studies point to different quantum resources, such as entanglement and non-locality [7], quantum coherence [8] and contextuality [9]. But none of these quantum features seems to suffice on its own to explain quantum advantage, due to results such as the Gottesman-Knill theorem [10] (though see for example the discussion in [6]). Furthermore, calibration and benchmarking quantum computers is usually a complicated business. As extant quantum processors are very sensitive to noise, it is important to develop different tools to assess their capabilities (see, for example, [11, 12, 13, 14]). The general question in which we are interested is: how are quantum resources distributed among the states that a concrete quantum device can produce? In this work we propose a method to assess that.
We study the resources produced in different architectures of quantum computers using states generated by quantum random circuits (QRC) [15]. We focus on non-locality by using the Mermin [16] and the Svetlichny inequalities [17, 18]. For completeness, we compare non-locality with entanglement by appealing to different measures, such as tangle [19] and entanglement for multi-partite systems [20]. We also analyze a quantifier of how much non-Clifford gates are present in a given circuit which is known as quantum magic [21].
@@
We first perform a theoretical study based on numerical simulations using the Amazon Braket SDK [22]. We analyze the performance of the devices with different levels of noise and number of shots, and compare universal (Clifford ++ T) vs non-universal (Clifford) sets of elementary gates. Given a set of physical features which are thought of as a resource for quantum computing (e.g., non-locality), we quantify up to which extent a device can give place to a state space rich enough to have the desired resources. The most salient difference between the universal and non-universal sets of gates analyzed in this work, is the discreteness in the possible values of the resources distribution of the latter. Indeed, while a device using only the Clifford set of gates can produce states with maximal entanglement and non-locality, the possible values of these magnitudes are highly concentrated in a few discrete values. We also show how the resources distribution is deteriorated with noise for architectures based on universal sets of gates.
Our findings are also relevant for the certification of quantum devices. By quantifying to which extent the resources distribution departs from the ideal case, one can have a global evaluation of the effective state space that a concrete quantum processing unit (QPU) can generate. The proposed methodology is different from a test merely based on preparing a concrete family of states, such as generalizations of the Greenberger–Horne–Zeilinger (GHZ) state, or testing a particular algorithm. Our developments allow to assess the capabilities of a quantum device as a whole, by using one of the main features of fundamental physics, namely, non-locality (see also [23, 24, 25] for other studies based on nonclassical correlations in relation to quantum processors). Our methodology can also be used to test the performance of a specific set of target qubits, a task that is relevant for calibration purposes.
Following our numerical simulations, we have also performed an implementation on quantum processing units based on Trapped Ions [26] and Superconducting Qubits [27]. More specifically, we use IonQ [28] and IQM [29] quantum processors. The experimental results show that our approach can be used as an effective tool to assess whether a given QPU satisfies the end user’s needs, since it yields very precise information about what portion of the state space can be reached as the number of qubits grows. These experiments also raise the question about to which extent one can affirm that a given QPU is capable of simulating non-local states.
The paper is organized as follows. First we give a brief overview of the Svetlichny and Mermin inequalities for multiple qubits and the notion of QRC. Next, we show the results of simulations of the violation of Mermin’s and Svetlichny’s inequalities for different numbers of qubits, shots, and noise levels, and compare the results of universal vs non-universal sets of quantum gates. We also include the analysis of different entanglement measures and quantum magic. We then propose a certification protocol based on our theoretical developments, and perform some experiments on IonQ and IQM processors to show how our approach can be useful to assess the performance of different architectures. In the Discussion section, after making some remarks about the interpretation of the experimental results, we draw our conclusions and leave open some questions for future work.
II Detecting non-locality
Figure 1: Schematic representation of IQM Garnet’s connectivity, where the three qubits Svetlichny inequality is to be tested on qubits 11, 22, and 44. Each arrow represents a possible spin measurement on the chosen qubits. There are two directions per qubit, described by blue or yellow arrows. Thus, there are eight correlation terms in total (see Eqs. (9) and (10) in Appendix A), which are schematically depicted as eight orientation choices on the right box.
Since the celebrated work of John S. Bell [30], the notion of quantum non-locality entered into the stage of quantum physics and gave place to heated foundational debates. This phenomenon, understood as a violation of local realism, is mathematically characterized by the lack of a probability distribution that simultaneously reproduces the observed correlations and meets a certain separability condition [31]. The failure of local-realism is expressed in the violation of a family of inequalities involving correlations among the different parties of the system.
For multi-component quantum systems, different types of non-locality inequalities exist. This work focuses on the Mermin [16] and Svetlichny [17, 18] inequalities (see Appendix A for details). A key difference is that violating a Svetlichny inequality implies the system exhibits genuine multipartite non-locality. This means that there exists no bipartition of the system for which the observed correlations can be explained by any local-realistic model. In this sense, a violation of the Svetlichny inequality allows to discard what is known as hybrid local realism.